OpenStudy (anonymous):

Tyler is constructing the frame for a triangular sand box. He has cut two pieces of wood that measure 8 ft and 4 ft. Which of the following could be the length of the third piece of wood that he needs to cut? 14 3 6 12

OpenStudy (kropot72):

Either 6 feet or 12 feet will work..

OpenStudy (anonymous):

thanks but it can only be one

OpenStudy (kropot72):

6 feet will make all three angles acute. So it gets my vote :)

OpenStudy (anonymous):

thank you !

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

In order to solve this question, you will need to use the Triangle inequality. It states that for any triangle, the sum of the lengths of any two sides must be greater than the length of the remaining side. If this triangle has 3 sides X = 8, Y = 4 and Z = ? then we have: Z < X + Y = 8 + 4 = 12 => Z < 12 We should apply the same rule for other sides as well. We have: X < Z + Y => 8 < Z + 4 => 4 < Z Therefore 4 < Z < 12 The only possible answer is 6.

OpenStudy (anonymous):

Here's the general form of the Triangle Inequality: For a triangle with 3 sides X, Y and Z we have: X < Y + Z Y < X + Z Z < X + Y

OpenStudy (anonymous):

thanks Parsa for explaining everything il have to redo this whole quiz and this will help me thanks i owe u a medal

OpenStudy (anonymous):

Haha, it's ok.